Circular No. 7813

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002aq IN MCG –01-7-35 P. Martin and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova (mag ∼ 18.7) on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Feb. 1.2 and 2.2 UT. SN 2002aq is located at α = 2^{h}41^{m}57^{s}.44, δ = −6^{o}47′46′′.4 (equinox 2000.0), which is 7′′.3 east and 14′′.2 south of the nucleus of MCG –01-7-35. A KAIT image taken on Jan. 20.2 already showed a hint of the new object, while an image taken on Jan. 11.2 showed nothing at this position (limiting mag ∼ 19.5).

COMET C/2002 C1 (IKEYA-ZHANG) Precise astrometry (Feb. 1–2) and the preliminary parabolic orbital

elements given below appear on MPEC 2002-C03. Visual m_{1} and comadiameter estimates: Feb. 1.910 UT, 7.5:, 5′ (P. M. Raymundo, Salvador, Brazil, 0.25-m reflector; independent discovery); 2.081, 9.5:, 3′ (A. Hale, Cloudcroft, NM, 0.41-m reflector; thin clouds); 2.43, 8.8, ∼ 4′ (K. Yoshimoto, Yamaguchi, Japan, 20×100 binoculars); 2.47, 8.8, 3′ (D. Zhang, Kaifeng, Henan, China, 0.20-m f/4.4 reflector, 28×, as used for the discovery on Feb. 1.47; comet more condensed than on previous day); 2.53, 9.1, 3′ (A. Pearce, Nedlands, W. Australia, 0.2-m reflector); 2.53, 8.5, 5′ (N. Brown, Quinns Rocks, W. Australia, 0.15-m refractor).

T = 2002 Mar. 8.912 TT  ω = 19.^{o}237
Ω = 111.715 }2000.0
q = 0.49127 AU  i = 26.341
2002TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Jan. 26 23^{h}58.^{m}97 −20^{o}32.^{′}2 1.410 1.055 48.^{o}4 44.^{o}2 9.5
28 0 01.99 −19 40.4 1.386 1.020 47.3 45.2 9.3
30 0 05.11 −18 46.6 1.361 0.986 46.3 46.3 9.1
Feb. 1 0 08.31 −17 50.4 1.335 0.951 45.4 47.5 8.9
3 0 11.58 −16 51.8 1.307 0.916 44.4 48.8 8.7
5 0 14.93 −15 50.5 1.279 0.882 43.5 50.3 8.5
7 0 18.33 −14 46.2 1.249 0.847 42.5 51.9 8.3
9 0 21.77 −13 38.8 1.218 0.813 41.6 53.7 8.0
11 0 25.24 −12 27.7 1.186 0.779 40.7 55.7 7.8
13 0 28.71 −11 12.8 1.153 0.746 39.9 58.0 7.5
15 0 32.14 − 9 53.6 1.119 0.714 39.0 60.5 7.3
2002 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green